Admission at the Computer Expo is $1.50 for kids and $4.00 for adults. On day, 2200 people went to the Expo and $5050 was collected. How many kids and how many adults attended that day?
Answer by mananth(16946)

x Kids
y Adults
Total attended
x+y = 2200 .............1
Kids 1.5 Adults 4 Amount collected 5050
1.50 x + 4.00 y = 5050 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 1
-4 x -4 y = -8800
1.50 x + 4 y = 5050
Add the two equations
-2.50 x = -3750
/ -2.50
x = 1500.00
plug value of x in (1)
1.00 x + 1.00 y = 2200.00
1500.00 + 1.00 y = 2200.00
1.00 y = 2200.00 + -1500.00
1.00 y = 700.00
y = 700.00
1500 Kids
700 Adults
m.ananth@hotmail.ca